HOW MUCH analyte is needed in case you are given 500, 250, or 200 mL volumetric flask to make your stock solutions.

maroon stock solution: Maroon

MW=  496.43 g/mol, Density:= 4.9707 g/L

gold stock solution: Gold

MW=  534.36 g/mol, Density=  5.1050 g/L
Answer #1

density = mass/volume

for maroon solution -- 4.9707= weight/1000ml

that means amount of analyte in 1000ml= 4.9707g

in 500ml it will be= 4.9707÷2

= 2.48535g

in 250 ml it will be =4.9707÷4

=1.24267g

for gold solution amount of analyte in 1000 ml = 5.1050g

hence in 500ml its amount will be = 5.1050÷2

=2.5525g

in 250 ml its amount will be needed= 5.1050÷4

=1.27625g